OpenStudy (anonymous):

Solve: x2 – 16x + 15 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two rots of x are: \(\frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

x^2 or 2X?

OpenStudy (anonymous):

(x-15)(x-1) Is more easy to see, 2 numbers that a.b = 15 and plus = -16 (-15)(-1) = 15 -15-1 = -16